OpenStudy (anonymous):

What is the slope of the line passing through (-3, 5) and (5, -3)?

jimthompson5910 (jim_thompson5910):

Subtract the y coordinates (first from the second) -3 - 5 = -8 Then subtract the x coordinates (same order) 5 - (-3) = 5+3 = 8 Now divide the two differences (y difference over x difference) -8/8 = -1 So the slope is -1

OpenStudy (anonymous):

Thank's so much for the help! :)

jimthompson5910 (jim_thompson5910):

you're welcome